Critical angle of glass is  θ1 and the critical angle of water is θ2. Calculate the critical angle for water and glass surface. (μg = 3/2, μw = 4/3)


Grade:10

2 Answers

Amit Askiitiansexpert
13 years ago

Dear kaushik

The critical angle is the angle of incidence above which total internal reflection occurs. The angle of incidence is measured with respect to the normal at the refractive boundary. The critical angle θc is given by:

\theta_c = \arcsin \left( \frac{n_2}{n_1} \right),
critical angle = arcsin(4/3 / 3/2) = 62.73 degres


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed  solution very  quickly.



All the best.

 

Regards

Amit - IT BHU




Kaushik Ram
13 years ago

Dear me The critical angle is the angle of incidence above which total internal reflection occurs. The angle of incidence is measured with respect to the normal at the refractive boundary. The critical angle θc is given by: critical angle = arcsin(4/3 / 3/2) = 62.73 degres